OpenStudy (anonymous):

Perform indicated operation and, if possible, simplify. Please show all work. -24/55/39/40

OpenStudy (lilg132):

-64/143

OpenStudy (anonymous):

how did you get this

OpenStudy (lilg132):

-24 - 40 = -64 55 + 39 + 40 = 143 = -64/143

OpenStudy (anonymous):

thank you